40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu May 14 17:41:10 2015, rana, Update, PEM, weather station and Guralp maintenance GurPost_150514.pngsecretWeatherTrends.png
    Reply  Sat May 16 21:40:29 2015, rana, Update, PEM, Guralp maintenance GurPost_150516.png
       Reply  Thu May 28 10:42:19 2015, Steve, Update, PEM, bad Guralp X-cable Gurs180dXbad.pngswappedGurs.png
          Reply  Thu May 28 15:10:44 2015, rana, Update, PEM, Seismic Confusion 
    Reply  Tue May 26 11:05:10 2015, Steve, Update, PEM, worked around ETMY seism. ETMYseismic.pngcut.jpg
Message ID: 11295     Entry time: Sat May 16 21:40:29 2015     In reply to: 11291     Reply to this: 11329
Author: rana 
Type: Update 
Category: PEM 
Subject: Guralp maintenance 

Tried swapping cables at the Guralp interface box side. It seems that all of our seismic signal problems have to do with the GUR2 cable being flaky (not surprising since it looks like it was patched with Orange Electrical tape!! rather than proper mechanical strain relief).

After swapping the cables today, the GUR2 DAQ channels all look fine: i.e. GUR1 (the one at the Y end) is fine, as is its cable and the GUR2 analog channels inside the interface box.

OTOH, the GUR1 DAQ channels (which have GUR2 (EX) connected into it) are too small by a factor of ~1000. Seems like that end of the cable will need to be remade. Luckily Jenne is still around this week and can point us to the pinout / instructions. Looks like there could be some shorting inside the backshell, so I've left it disconnected rather than risk damaging the seismometer. We should get a GUR1 style backshell to remake this cable. It might also be possible that the end at the seismometer is bad - Steve was supposed to swap the screws on the granite-aluminum plate on Thursday; I'll double check.

Attachment 1: GurPost_150516.png  50 kB  Uploaded Sat May 16 23:37:05 2015  | Hide | Hide all
GurPost_150516.png
ELOG V3.1.3-